OpenStudy (anonymous):

How to find the close form of \(a_n =\frac{1}{8}a_{n-2} -\frac{1}{4}a_{n−1}\) given that \(a_1 = 0, a_2 = -1\)?

OpenStudy (ikram002p):

what is a_0 and a_1 ? or any initial a's ?

OpenStudy (anonymous):

a_0 = 0, a_1 = -1

OpenStudy (ikram002p):

ok , expand only unfortunately I have to go cuz I have a class brb

hartnn (hartnn):

get a_2 , a_3 ... and search for pattern... I'll think about an alternate approach..

OpenStudy (anonymous):

Edit: a1 = 0, a2 = -1, a3 = 1/4, a4 = -3/16

OpenStudy (anonymous):

I would like to know if there is a systematic way to solve such difference equation.

hartnn (hartnn):

are you expecting something like this : \(\Large a_n = (-1/4)a_{n-1} + (1/8-1/4)\sum \limits_{i=2}^{n-2}a_i+(1/8)a_1 \)

OpenStudy (anonymous):

Ehm.. Not really...

hartnn (hartnn):

ok, in terms of n only

OpenStudy (anonymous):

Yup!

OpenStudy (experimentx):

sorry .. i forgot to reply. do you know z-transform?

OpenStudy (anonymous):

I learnt it in EEE 2 years ago, but forgot most of it

OpenStudy (experimentx):

there are two ways of doing it .. actually three. #1 find out the general pattern as you list the terms. #2 use z-transform #3 FInd the eiven vectors and the eigen values of the equation